Presentation is loading. Please wait.

Presentation is loading. Please wait.

Year End Technical Stop DT work Mu Barrel Workshop 13 October 2011 A. Benvenuti (INFN Bologna)

Similar presentations

Presentation on theme: "Year End Technical Stop DT work Mu Barrel Workshop 13 October 2011 A. Benvenuti (INFN Bologna)"— Presentation transcript:

1 Year End Technical Stop DT work Mu Barrel Workshop 13 October 2011 A. Benvenuti (INFN Bologna)

2 Year End Technical STOP Technical coordination “Retreat” in Divonne on September 28-29 to assess opportunity/necessity of opening the detector: Main driver being access to Tracker for studying the humidity control, in view of operating at low temperature, - 20C after LS1  Not cost effective within a 13 weeks YETS; additional 4 weeks needed.  Not ruled out since LHC plans not yet final (Thermal amplifier tests..) but unlikely unless start of LS1, 17 November 2012, also shifts  Not very interesting for DT work since, at best, only limited access would be possible to wheels

3 Nicola Bacchetta

4 Year End Technical STOP Provisional Schedule 7/12 @ 18:00 Beam OFF (year end calibrations) 8/12 @ 08:00 Magnet Ramp Down (DT shutdown ) 8/12 @ 14:00 Water Cooling Maintenance 12/12 AUG tests UXC closed 13/12 CASTOR Removal Test 21/12 HF on the floor 09/01 Cooling is back 06/02 Skeleton Shift crew 24/7 08/02 first MWGR 23/02 Magnet Ramp to full field (Cosmics data taking) 07/03 Beams back in LHC + 3 weeks commissioning with beams

5 HV Problems Summary Compared to 9 chambers with HV problems and 132 channels lost at the end of 2010 (YB-2 MB4 S07 is intermittent, not included in 2010 statistics) September 28 2011

6 A Snapshot of DT and RPC (Barrel) Maintenance Work (as of today) SUMMARY Gianni Masetti: DT dead channels grand total = 1050 (0.6%) Only work requiring access to the detector is entered in the table Only RPC maintenance that requires moving the chambers is included September 28 2011

7 7 1 4 2 3 5 6 7 8 9 10 11 12 1 4 2 3 5 6 7 8 9 10 11 12 1 4 2 3 5 6 7 8 9 10 11 12 1 4 2 3 5 6 7 8 9 10 11 12 1 4 2 3 5 6 7 8 9 10 11 12 DT HV Phi +2 +1 0 -2 DT HV Theta RPC Extract RPC Replace Problems in DT or RPC Requires uncabling

8 Can DTs Benefit from Opening the Detector ? It depends for how long and how far: 1) Opening up to YB±2 allows access to the MiniCrates and inspection of cooling and gas connections 2) Opening up to YB±1 allows HV repairs on Phi SL of two chambers in YB-2 and one in YB+2 plus maintenance of YB±1 3) Opening up to YB±1 allows also access to HO RBX in the movable wheels. However, for YB±1 this requires removing DT gas and cooling lines from MB1. 4)Chamber extraction out of the iron pockets is not feasible in any “normal” scenario for the end of the year TS

9 Caveats Opening and closing the detector are delicate operations that might damage some components. The risks must be weighted against the gains from the possible fixes. Each DT intervention requires time consuming tests with cosmics data taking that are possible only when the general infrastructures (power, cooling, gas and DAQ) are operational The wheels cannot be closed until the tests are successfully completed  For the DTs, the balance is not in favor of opening the detector

10 DT Off Detector Maintenance: LV Connectors Melted connector found on 28/ 10/ 2009 Weak point identified in the poor match of cable (Powerclaw) and CAEN supply connectors (PP75) exacerbated by rigid cables All jumper cables and connectors were replaced using very flexible cable in the 2009 -10 Technical Stop Monitoring of temperatures and Vdrops was extended considerably Still several problems in 2010 and 2011 Major cause of UXC access requests for DT The way it should be

11 A Radical Solution “Pigtails” with AP connectors soldered on the LV modules 4 modules have been modified and are in use since the July TS Some numbers: 26 A3050 and 2 A3100/ wheel 112 pigtails/wheel 1 – 1.5 hours/ module Modules test : 8-12/day  1 Wheel/ week All pigtails should be ready before starting 2 Soldering stations Soldering procedures and trained personnel (developed by M. Pegoraro) Best place is in S1 (proximity to detector)  Decision on the modules to modify, types and quantity, by November at latest

12 LV Pigtails (continued) Last Friday, Cristina presented the plan and requests for this intervention during the TC meeting Well received, HCAL interested in applying this solution Space identified in area between S1 and elevator => OK Draft of “Procedures Agreement” prepared by Cristina  Critical items: Test system(s) (Daniele F.) Components: cable, connectors, PENETROX (3 bottles from MJ)… Derogations Pigtails preparation ( 26 A3050, 2 A3100, 8MAO/YB) = 720 pigtails in total (5 wheels) plus spares Decouple RP clearance from intervention Logistics (moving in and out the modules), dressing cables in tower racks and testing afterwards is a very demanding task  Manpower requirements: 2 +1 technicians for soldering, 1+1 for moving and cabling  Total of 4-5 Technicians for 5 weeks + physicists for testing

13 LV Connectors (The Challenge) Define pigtails lengths and strain relief during next TS Identify equipment and procedures for moving modules MAO extraction and transportation is a major challenge (pain)  130 A3050, 10 A3100, 40 MAO

14 YETS 2011 (In Addition) Check A3009 connections (Hotspots and currents to be done during next TS, magnet OFF, not enough time before YETS power off) Improve TJB noise shielding (present aluminum foil and aluminum tape turkey wrap is not viable for the long run) Insert new TJBs and check persistence of HV problems Gas system maintenance/improvements to be agreed with Roberto Guida: - Calibrate ~ 10 flow-meters - Replace humidity analysis system in four wheels - Install and commission the “inline gas chromatograph” - Isolate the VDC lines from the O2 reference gas - Determine time required to clear out gas in VDC during analysis cycle (now set to 1 hour) Some minor VDC interventions: new gas line connections, VDCs intercalibration, improved air tightness by adding a VDC “hood”  AOB, depending on what happens between now and 06 December

15 First measurement after the “reference gas” point is out of norm for all VDC. Effect disappears if measurement time is increased to ~4 hours Measurements for YB±1 have ~ factor 2 larger errors  Daniel is investigating the cause

16 Summary The LV supplies modification is the main activity for this YETS All pigtails must be ready before January 16 for the intervention to succeed  The intervention can proceed in stages: modify all MAO, all A3100 (SC), all A3050 in a wheel in order to minimize the risks In case the detector will be opened we can take the opportunity to fix the know MC problems in YB ± 2 and inspect gas and water connections. This would not interfere with the other DT planned activities  But of course, plans can change…….

17 Backup

18 Extraction tool mounted on MB2S09. Similar position as needed for YB0 MB3S06 and MB2S09. The SL cover to open is located in the upper part of the chamber for MB2S09 and in the bottom part for MB3S06.

19 Interventions on top sectors are done with heavy equipment above the beampipe

20 And, to a lesser extent, also work on YB+1

21 1)Remove MABs: S03/S04, S01/S02 and S05/S06 (Maf + alignment team) 2)Remove covers from radial cable trays 3)Remove Thermal Shields from Vacuum Tank to access ECAL LV Patch Panels, 2 TS/sector 4)Disconnect ECAL LV Cables ( 98 ) from PP on VT 5)Pull cables to Sector periphery and secure on cable trays ( 2 persons on scissor lift + 1 person on periphery) 6)Mount platforms on VT, where needed 7)Remove ECAL radial cable trays 8)Disconnect, pull-out and secure the optical fibers (DT technicians) 9)Disconnect DT and RPC cables ( DT+RPC tecs) 10)Pull cables to the periphery and secure them 11)Disconnect Gas and cooling (DT+RPC techs) Z0+ Cables Removal Platforms are not needed for the cables removal in S06. The extraction tool can be mounted with a scissor lift The HV repair can be carried out from the scissor lift

Download ppt "Year End Technical Stop DT work Mu Barrel Workshop 13 October 2011 A. Benvenuti (INFN Bologna)"

Similar presentations

Ads by Google